### 03 Apr 2026 v2

# Version-Robust Methods for Identifying Minimal Sufficient Statistics

Abstract
Let $f_{\theta}$ be the joint density of a random sample $X$. A frequently used criterion asserts that a statistic $T(X)$ is minimal sufficient if, for any sample points $x$ and $y$, $T(x)=T(y)$ exactly when there exists a finite constant $h_{xy}>0$, independent of $\theta$, such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for all $\theta$. We show that this criterion is false in general via a counterexample exploiting the non-uniqueness of versions of Radon–Nikodym derivatives. Although Sato (1996) established sufficient regularity conditions for the validity of this criterion, these conditions are frequently intractable to verify in practice. We resolve this limitation by introducing a version-robust method applicable whenever sufficiency is known. Moreover, our method allows us to generalize Sato’s approach from Euclidean settings to arbitrary analytic Borel sample spaces and separable measurable statistic spaces. We also obtain a method for exponential-family densities under verifiable hypotheses. Taken together, these results clarify when pointwise likelihood-ratio arguments for minimal sufficiency are mathematically sound in irregular settings. Finally, we construct a counterexample demonstrating that a distinct criterion for minimal sufficiency due to Pfanzagl (1994, 2017) similarly fails in the absence of supplementary hypotheses. Identifying minimal sufficient statistics is important not only for parsimonious data reduction but also because, in models admitting complete sufficiency, such statistics provide a practical route to the complete sufficient structure underlying optimal estimation and prediction.
Keywords: Analytic Borel spaces; exponential families; irregular models; minimal sufficiency; minimal sufficient statistics; Pfanzagl’s criterion; Radon–Nikodym derivatives; Sato’s method.
1 Introduction
Minimal sufficiency is a data-reduction problem in dominated statistical models: one seeks a statistic that retains all inferentially relevant information in the sample while being minimal among sufficient reductions. A common strategy is to compare sample points through proportionality relations between likelihoods or densities. According to the Lehmann–Scheffé theorem (Liese and Miescke, 2008, Theorem 7.17), one route to finding a uniformly minimum-variance unbiased estimator is to condition an unbiased estimator on a statistic that is both sufficient and complete. In practice, however, complete sufficient statistics are often difficult to construct directly. Nevertheless, the relation between completeness and minimal sufficiency remains important: whenever a model admits a sufficient and complete statistic, every minimal sufficient statistic is automatically complete. For this reason, methods for identifying minimal sufficient statistics are useful not only for data reduction itself but also, in such models, as an indirect route to complete sufficiency in applications such as estimation theory. This connection is also relevant for prediction, since in such models an identified minimal sufficient statistic may serve as the complete sufficient statistic underlying the Bayes-optimal prediction procedures with frequentist coverage control developed in Hoff (2023).
Let $\Theta$ be the parameter space and $T$ be a statistic defined on the statistical model. Informally, $T$ is said to be sufficient if, and only if, the data distribution conditional on $T$ does not depend on $\theta\in\Theta$, and $T$ is said to be minimal sufficient if, and only if, it is sufficient and, given any sufficient statistic $S$, there exists a measurable map $f$ such that $T=f\circ S$ almost everywhere. Throughout this paper, the formal definitions of statistical models, statistics, sufficiency and minimal sufficiency, standard and analytic Borel spaces, countably generated spaces and separable measurable spaces are collected in Appendix A.
In this paper, we discuss the limitations of two pointwise criteria for identifying minimal sufficiency, propose version-robust methods that avoid such limitations, and also generalize a method proposed by Sato (1996). The first and widely cited criterion considered in this paper can be found in texts such as Schervish (1995), see his Theorem 2.29, Wasserman (2004), see his Theorem 9.36, and Mavrakakis and Penzer (2021), see their Proposition 10.1.13, and it is stated as follows.
Criterion $1.$
Let $f_{\theta}(x)$ be a joint density of a random sample $X$. A statistic $T(X)$ is minimal sufficient if, for any two sample points $x$ and $y$, we have $T(x)=T(y)$ iff there exists a finite constant $h_{xy}>0$ independent of $\theta$ such that
$f_{\theta}(y)=f_{\theta}(x)h_{xy}\quad\mbox{for every}\ \theta\in\Theta.$
In this formulation, the restriction applies over the entire (possibly uncountable) parameter space, which makes the pointwise proportionality relation sensitive to the chosen versions of the densities and allows us to develop Counterexample 2.1. Although widely quoted as a method for minimal sufficiency (see also Young and Smith, 2005; Makarov and Podkorytov, 2013; Olive, 2014; Gasperoni et al., 2025), Criterion 1.1 is often stated without explicit regularity hypotheses and is false in full generality. In particular, since densities are only defined almost everywhere, one can modify versions on null sets in a $\theta$-dependent way and thereby change the pointwise proportionality relation.
The formal conditions under which a reformulation of this method holds, preventing our counterexample, were first established by Sato (1996). This method is restricted to Euclidean spaces and is not simple to apply in practice. To bypass this practical difficulty, when a candidate statistic is already known to be sufficient (a condition that is typically easy to verify via the Neyman–Fisher factorization theorem), we introduce a more direct version-robust minimality method (Method 3.1). We then use this approach to extend Sato’s method beyond Euclidean spaces to analytic Borel sample spaces and separable measurable statistic spaces (Method 3.2), and to establish a minimal-sufficiency method for exponential-family densities (Method 3.3).
The second criterion considered in this paper was developed by Pfanzagl (1994) and reformulated in Pfanzagl (2017). This approach is formally stated in terms of dominated statistical models, countably generated measurable spaces, measurable functions and standard Borel spaces, which are revisited in our Appendix A. By using these concepts, we present Pfanzagl’s criterion.
Criterion 1.2.
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$ (i.e., $P_{\theta}$ is absolutely continuous w.r.t. $\mu$ for every $\theta\in\Theta$). Suppose that $(\mathcal{X},\Sigma)$ is countably generated. Let $(\mathcal{T},\Sigma_{\mathcal{T}})$ be a standard Borel space and $T:\mathcal{X}\to\mathcal{T}$ a measurable function. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Suppose that, for each $\theta\in\Theta$, we have $f_{\theta}=g_{\theta}(T)h$, where $h:\mathcal{X}\to\mathbb{R}$ and $g_{\theta}:\mathcal{T}\to\mathbb{R}$ are two non-negative measurable functions. If there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that, for any $t_{1},t_{2}\in\mathcal{T}$ satisfying $g_{\theta}(t_{1})=g_{\theta}(t_{2})$ for every $\theta\in\Theta_{0}$, we have $t_{1}=t_{2}$, then $T$ is minimal sufficient.
This criterion corresponds to Theorem 1.4.4 of Pfanzagl (1994). A later version of this result in Pfanzagl (2017) is stated with the additional requirement that $\{P_{\theta}\}_{\theta\in\Theta_{0}}$ is dense in $\{P_{\theta}\}_{\theta\in\Theta}$ w.r.t. the total variation distance, defined as $d(P,Q):=\sup_{E\in\Sigma}|P(E)-Q(E)|$. However, as noted at the beginning of the proof in Pfanzagl (1994), this density assumption is not necessary and may be omitted. Note also that Pfanzagl’s criterion imposes a restriction only on at most a countable subset of the parameter space. However, it still does not hold without additional assumptions, as shown by our Counterexample $2.$
The remainder of this paper is organized as follows. In Section 2, the limitations of these methods are demonstrated through specific counterexamples that motivate the need for version-robust methods. Section 3 introduces the corrected and generalized methods that form the core contribution of this work, followed by examples illustrating their applications. Section 4 is dedicated to the formal proofs of these methods. Appendix A presents the measurable-space background used throughout the paper, namely, statistical models (Definition A.1), sufficient statistics (Definition A.2), standard Borel spaces (Definition A.3), countably generated spaces (Definition A.4), separable measurable spaces (Definition A.5), and analytic Borel spaces (Definition A.6). We denote the set of natural numbers by $\mathbb{N}=\{0,1,2,\ldots\}$. Moreover, we will abbreviate the expressions “if, and only if,” and “with respect to” by “iff” and “w.r.t.”, respectively.
2 Counterexamples
In this section, we construct counterexamples showing Criteria 1.1 and 1.2 require further regularity hypotheses. The first, similar to the example from Barndorff-Nielsen et al. (1976) and closely related to Example 1 of Taraldsen (2026), perturbs the Gaussian density at a single $\theta$-dependent point, falsifying Criterion $1.$ The second invalidates Criterion 1.2 on a finite probability space. These counterexamples motivate the methods introduced in Section 3.
Counterexample 2.1.
Let $X_{1},\cdots,X_{n}$ be a random sample with normal distribution $N(\theta,1)$, $\theta\in\mathbb{R}$, with $n\geq 2$. Using the Neyman-Fisher Factorization Theorem, we can conclude that $\sum_{i=1}^{n}X_{i}$ is sufficient. Denote by $\lambda^{n}$ the Lebesgue measure on $\mathbb{R}^{n}$. Let $g:\mathbb{R}\to\mathbb{R}^{n}$ be any surjective function. Define $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ by
$f_{\theta}(x_{1},\cdots,x_{n}):=\frac{1}{(2\pi)^{n/2}}\exp\left\{-\frac{1}{2}\left(\sum_{i=1}^{n}x_{i}^{2}\right)+\theta\left(\sum_{i=1}^{n}x_{i}\right)-\frac{n}{2}\theta^{2}\right\}\mathbf{1}_{\mathbb{R}^{n}\setminus\{g(\theta)\}}(x_{1},\cdots,x_{n}),$
where $\mathbf{1}_{\mathbb{R}^{n}\setminus\{g(\theta)\}}$ denotes the indicator function of the set $\mathbb{R}^{n}\setminus\{g(\theta)\}$. Since $\lambda^{n}(\{g(\theta)\})=0$, $f_{\theta}$ is a joint density of $X:=(X_{1},\cdots,X_{n})$. Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the identity function. Since $g$ is surjective, we infer that, for any $x,y\in\mathbb{R}^{n}$, we have $T(x)=T(y)$ iff there exists a finite constant $h_{xy}>0$ independent of $\theta$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\mathbb{R}$. (Indeed, if $x\neq y$, we can choose a $\theta\in\mathbb{R}$ such that $g(\theta)=y$ due to surjectivity. For this $\theta$, the indicator function in the definition of $f_{\theta}$ causes $f_{\theta}(y)=0$, while $f_{\theta}(x)\neq 0$, making the equality $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ impossible for any $h_{xy}>0$.) Hence, according to Criterion 1.1, $X=T(X)$ is a minimal sufficient statistic, which implies, recalling that $\sum_{i=1}^{n}X_{i}$ is sufficient, that there exists a measurable function $u:\mathbb{R}\to\mathbb{R}^{n}$ such that $(x_{1},\cdots,x_{n})=u(\sum_{i=1}^{n}x_{i})$ $P_{\theta}$-a.e. for every $\theta\in\mathbb{R}$, where $P_{\theta}$ is the probability measure given by $P_{\theta}(E):=\int_{E}f_{\theta}d\lambda^{n}$. Since $f_{\theta}>0$ $\lambda^{n}$-a.e. for every $\theta\in\mathbb{R}$, we conclude that $(x_{1},\cdots,x_{n})=u(\sum_{i=1}^{n}x_{i})$ $\lambda^{n}$-a.e., a contradiction. To see this, define $A:=\{x\in\mathbb{R}^{n}:u(\sum_{i=1}^{n}x_{i})=x\}$ and $V:=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$. It is straightforward to prove that AA is measurable and $(A-A)\cap(V\setminus\{0\})=\emptyset$, in which $A-A:=\{a_{1}-a_{2}:a_{1},a_{2}\in A\}$. However, if $\lambda^{n}(A)>0$, then, by Steinhaus Theorem, we know that AAA-A contains an open ball centered at 0, allowing us to conclude that $(A-A)\cap(V\setminus\{0\})\neq\emptyset$.
Observe that, for each $\theta$, the function $f_{\theta}$ differs from the usual $N(\theta,1)^{\otimes n}$ joint density only on the $\lambda^{n}$-null set $\{g(\theta)\}$. Hence the induced measure $P_{\theta}(E):=\int_{E}f_{\theta}\,d\lambda^{n}$ coincides with the usual $N(\theta,1)^{\otimes n}$ law. The counterexample therefore exploits the non-uniqueness of Radon–Nikodym derivatives (choice of version), not a change in the statistical model.
The previous counterexample demonstrates that Criterion 1.1, as commonly stated in the literature, does not hold in general. The proof that often accompanies this criterion, while seemingly correct, typically contains a subtle error. The flaw lies in an imprecise application of the Neyman-Fisher Factorization Theorem. Specifically, this factorization theorem states that if $f_{\theta}$ is a joint density of a random sample $X$ and $T(X)$ is a sufficient statistic, then there exist measurable functions $g_{\theta}$ and $h$ such that $f_{\theta}=g_{\theta}(T)h$ almost everywhere. The theorem, however, does not guarantee that this equality holds for every point. Therefore, for a pre-specified density $f_{\theta}$, a specific sample point $x$, and sufficient statistic $T$, one cannot guarantee that there exist measurable functions $g_{\theta}$ and $h$ such that $f_{\theta}(x)=g_{\theta}(T(x))h(x)$.
We conclude this discussion with the following counterexample, which shows that Criterion 1.2 does not hold without additional conditions.
Counterexample $2.$
[…]
We have $p_{\theta}(x)=g_{\theta}(T(x))h(x)$ for any $\theta\in(0,1)$ and $x\in\mathcal{X}$, where $h:\mathcal{X}\to\mathbb{R}$ is given by $h(1)=h(3)=1/3$ and $h(2)=h(4)=2/3$ and, for each $\theta\in(0,1)$, $g_{\theta}:\{0,1\}\to\mathbb{R}$ is given by $g_{\theta}(0):=1-\theta$ and $g_{\theta}(1):=\theta$. Note that $h$ and $g_{\theta}$ are non-negative measurable functions for every $\theta\in(0,1)$. Hence, using the Neyman-Fisher Factorization Theorem, we conclude that $T$ is sufficient.
Define $U:\mathcal{X}\to\mathcal{X}$ by $U(x):=x$. Then $U$ is measurable. We will now prove that $U$ is minimal sufficient using Pfanzagl’s method. We can write $p_{\theta}(x)=\tilde{g}_{\theta}(U(x))\tilde{h}(x)$ for any $\theta\in(0,1)$ and $x\in\mathcal{X}$, where $\tilde{h}:\mathcal{X}\to\mathbb{R}$ is given by $\tilde{h}(x):=1$, and $\tilde{g}_{\theta}:\mathcal{X}\to\mathbb{R}$ (for each $\theta\in(0,1)$) is given by $\tilde{g}_{\theta}:=p_{\theta}$. Note that $\tilde{h}$ and $\tilde{g}_{\theta}$ are non-negative measurable functions for every $\theta\in(0,1)$. It is straightforward to verify that, for any $t_{1},t_{2}\in\mathcal{X}$ satisfying $\tilde{g}_{\theta}(t_{1})=\tilde{g}_{\theta}(t_{2})$ for every $\theta\in(0,1)\cap\mathbb{Q}$, we have $t_{1}=t_{2}$. Therefore, as the set $(0,1)\cap\mathbb{Q}$ is countable, we conclude, using Pfanzagl’s method, that $U$ is minimal sufficient. Hence, since $T$ is sufficient, we conclude that there exists a measurable function $f:\{0,1\}\to\mathcal{X}$ such that for every $\theta\in(0,1)$, we have $U=f(T)$ $P_{\theta}$-a.e., implying that $U=f(T)$, because $P_{\theta}(\{i\})>0$ for any $\theta\in(0,1)$ and $i\in\mathcal{X}$.
Therefore, $1=U(1)=f(\mathbf{1}_{\{1,2\}}(1))=f(\mathbf{1}_{\{1,2\}}(2))=U(2)=2$, a contradiction.
The preceding counterexample shows that Theorem 1.4.4 in Pfanzagl (1994), which corresponds to Criterion 1.2, does not hold as stated. The gap in its proof lies in the application of an unjustified argument from Theorem 1.4.2 of that work. We now outline Pfanzagl’s argument to identify where the error occurs, adopting our notation instead of that of the original text.
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is countably generated. Let $(\mathcal{T},\Sigma_{\mathcal{T}})$ be a standard Borel space and $T:\mathcal{X}\to\mathcal{T}$ a measurable function. For each $\theta\in\Theta$, let $f_{\theta}^{\mu}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. We assume the following conditions: (a) for each $\theta\in\Theta$, we have $f_{\theta}^{\mu}=g_{\theta}^{\mu}(T)h^{\mu}$, where $h^{\mu}:\mathcal{X}\to\mathbb{R}$ and $g_{\theta}^{\mu}:\mathcal{T}\to\mathbb{R}$ are two non-negative measurable functions; and (b) there exists a subset $\{\theta_{n}\}_{n\in\mathbb{N}}\subseteq\Theta$ such that, for any $t_{1},t_{2}\in\mathcal{T}$ satisfying $g_{\theta_{n}}^{\mu}(t_{1})=g_{\theta_{n}}^{\mu}(t_{2})$ for every $n\in\mathbb{N}$, we have $t_{1}=t_{2}$.
Define $H^{\mu}:\mathcal{T}\to\mathbb{R}^{\mathbb{N}}$ by $H^{\mu}(t):=(g_{\theta_{n}}^{\mu}(t))_{n\in\mathbb{N}}$. In the proof of Theorem 1.4.4, Pfanzagl asserts that the statistic $T^{\mu}:\mathcal{X}\to\mathbb{R}^{\mathbb{N}}$ given by $T^{\mu}(x):=H^{\mu}(T(x))$ is minimal sufficient, invoking the proof of Theorem 1.4.2. However, the proof of Theorem 1.4.2 establishes only the existence of a probability measure $\mathbb{P}:\Sigma\to\mathbb{R}$ such that the statistic $T^{\mathbb{P}}:\mathcal{X}\to\mathbb{R}^{\mathbb{N}}$ given by $T^{\mathbb{P}}(x):=H^{\mathbb{P}}(T(x))$ is minimal sufficient, where $H^{\mathbb{P}}:\mathcal{T}\to\mathbb{R}^{\mathbb{N}}$ is defined by $H^{\mathbb{P}}(t):=(g_{\theta_{n}}^{\mathbb{P}}(t))_{n\in\mathbb{N}}$ with components $g_{\theta_{n}}^{\mathbb{P}}$ chosen such that the composition $g_{\theta_{n}}^{\mathbb{P}}(T)$ is a density of $P_{\theta_{n}}$ w.r.t. $\mathbb{P}$. Thus, the argument in Theorem 1.4.2 is purely existential; it constructs a specific collection of measurable functions yielding a minimal sufficient statistic, and does not imply that an arbitrary pre-specified collection, such as $\{g_{\theta_{n}}^{\mu}\}_{n\in\mathbb{N}}$, will produce a minimal sufficient statistic of the form $H(T(x))$. Moreover, Counterexample 2.2 demonstrates that this gap cannot be closed without further assumptions.
We conclude this section with a simple observation. Since the counting measure on a non-empty finite set, normalized by its cardinality, is a probability measure, we may multiply the densities in Counterexample 2.2 by 44; this shows that Criterion 1.2 remains false even if we additionally require $\mu$ to be a probability measure equivalent to $\{P_{\theta}\}_{\theta\in\Theta}$ (i.e., $\mu(E)=0$ if and only if $P_{\theta}(E)=0$ for every $\theta\in\Theta$).
3 Corrected and Generalized Methods
This section introduces version-robust and correct criteria for identifying minimal sufficient statistics, together with examples illustrating their application. The main idea is to replace pointwise likelihood comparisons over the whole parameter space by arguments based on countable subfamilies or verifiable approximation hypotheses. These criteria are rigorously proved in Section 4.
These methods are particularly useful in settings in which pointwise likelihood comparisons are delicate but still informative, such as symmetry models, models with parameter-dependent support, and models on analytic Borel sample spaces. In particular, Examples 3.13.4 illustrate Method 3.1, Example 3.6 illustrates Method 3.2, and Example 3.7 illustrates Method $3.$
Method 3.1.
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Let $(\mathcal{T},\Sigma_{\mathcal{T}})$ be a separable measurable space and $T:\mathcal{X}\to\mathcal{T}$ a measurable function. For any $x\in\mathcal{X}$ and $\Theta^{\prime}\subseteq\Theta$, define
$D(x,\Theta^{\prime}):=\big\{y\in\mathcal{X}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta^{\prime})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}.$
If $T$ is sufficient and there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that, for any $x,y\in\mathcal{X}$ satisfying $y\in D(x,\Theta_{0})$, we have $T(x)=T(y)$, then $T$ is minimal sufficient.
Counterexample 2.1 shows that if one permits $\Theta_{0}$ to be uncountable, then pointwise statements involving the densities $\{dP_{\theta}/d\mu\}_{\theta\in\Theta_{0}}$ become vulnerable to version dependence: one can choose versions in a $\theta$-dependent way on $\mu$-null sets and thereby alter the induced proportionality relation. In contrast, Method 3.1 avoids this pitfall by restricting to a countable subfamily $\Theta_{0}$, which enables the selection of versions that are consistent simultaneously for all $\theta\in\Theta_{0}$ outside a single $\mu$-null set.
The following examples illustrate the application of Method 3.1. The minimal sufficient statistic of Example 3.1 appears in Thomas and Anjana (2021).
Example 3.1.
Let $\Theta$ be the set of all probability densities $f:\mathbb{R}\to\mathbb{R}$ (w.r.t. the Lebesgue measure) such that f(x)=f(x)f(x)=f(-x) for every $x\in\mathbb{R}$. Let $X_{1},\cdots,X_{n}$ be a random sample with density $f\in\Theta$. Then $p_{f}:\mathbb{R}^{n}\to\mathbb{R}$ given by $p_{f}(x_{1},\cdots,x_{n}):=\prod_{i=1}^{n}f(x_{i})$ is a joint density of $(X_{1},\cdots,X_{n})$. Since $f$ is symmetric at 0 for every $f\in\Theta$, we have $p_{f}(x_{1},\cdots,x_{n})=\prod_{i=1}^{n}f(|x_{i}|)$ for every $f\in\Theta$, allowing us to conclude, using the Neyman-Fisher Factorization Theorem, that the statistic $(|X|_{(1)},\cdots,|X|_{(n)})$ is sufficient, where $|X|_{(i)}$ denotes the $i$-th order statistic of $(|X_{1}|,\cdots,|X_{n}|)$. Next, we will prove that $(|X|_{(1)},\cdots,|X|_{(n)})$ is also minimal.
For each $\alpha>0$, define $f_{\alpha}:\mathbb{R}\to\mathbb{R}$ by $f_{\alpha}(x):=\frac{1}{\pi\alpha(1+(x/\alpha)^{2})}$. That is, $f_{\alpha}$ is a density function of a random variable with distribution $\text{Cauchy}(0,\alpha)$. Hence, $f_{\alpha}\in\Theta$ for every $\alpha>0$. Define $\Theta_{0}:=\{f_{\alpha}:\alpha\in\mathbb{Q}_{+}\}$, where $\mathbb{Q}_{+}:=\mathbb{Q}\cap(0,\infty)$. Then $\Theta_{0}$ is a countable subset of $\Theta$. For every $x\in\mathbb{R}^{n}$, define $D(x,\Theta_{0}):=\big\{y\in\mathbb{R}^{n}:(\exists h_{xy}\in(0,\infty))(\forall f\in\Theta_{0})\big(p_{f}(y)=p_{f}(x)h_{xy}\big)\big\}$.
Let $x:=(x_{1},\cdots,x_{n}),y:=(y_{1},\cdots,y_{n})\in\mathbb{R}^{n}$ be arbitrary such that $y\in D(x,\Theta_{0})\,\,(i)$. Because of Method 3.1, to prove that $(|X|_{(1)},\cdots,|X|_{(n)})$ is minimal, it suffices to prove that $(|x|_{(1)},\cdots,|x|_{(n)})=(|y|_{(1)},\cdots,|y|_{(n)})$.
From $(i)$, we know that there exists $h_{xy}>0$ such that $p_{f_{\alpha}}(y)=p_{f_{\alpha}}(x)h_{xy}$ for every $\alpha\in\mathbb{Q}_{+}$. Hence, for each $\alpha\in\mathbb{Q}_{+}$, we have $\prod_{k=1}^{n}\frac{1}{\pi\alpha(1+(y_{k}/\alpha)^{2})}=h_{xy}\prod_{k=1}^{n}\frac{1}{\pi\alpha(1+(x_{k}/\alpha)^{2})}$, implying that $P(\alpha)=Q(\alpha)\,\,(ii)$ for every $\alpha\in\mathbb{Q}_{+}$, where $P,Q:\mathbb{R}\to\mathbb{R}$ are the polynomials given by $P(\alpha):=\prod_{k=1}^{n}(\alpha^{2}+x_{k}^{2})$ and $Q(\alpha):=h_{xy}\prod_{k=1}^{n}(\alpha^{2}+y_{k}^{2})$. Since $P,Q$ are even and continuous and $\mathbb{Q}_{+}$ is dense in $0,\infty)$, we conclude, using $(ii)$, that $P=Q$, allowing us to infer that $P$ and $Q$ have the same roots. Hence, $(|x|_{(1)},\cdots,|x|_{(n)})=(|y|_{(1)},\cdots,|y|_{(n)})$, since $\{\pm ix_{k}\}_{k=1}^{n}$ and $\{\pm iy_{k}\}_{k=1}^{n}$ are the sets that contain all roots of $P$ and $Q$, respectively, with $i$ denoting the complex number that satisfies $i^{2}=-1$.
Example 3.2.
Let $f:\mathbb{R}\to(0,\infty)$ be an integrable function. For each $\theta\in\mathbb{R}$, define $c(\theta):=1/\int_{\theta}^{\infty}f(x)dx$. Let $X_{1},\cdots,X_{n}$ be a random sample with density $p_{\theta}(x)=c(\theta)f(x)$, for $x>\theta$ and $p_{\theta}(x)=0$, otherwise. We will prove that the order statistic $X_{(1)}=\min(X_{1},\cdots,X_{n})$ is minimal sufficient. The function $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ given by
$f_{\theta}(x_{1},\cdots,x_{n}):=(c(\theta))^{n}f(x_{1})\cdots f(x_{n})\mathbf{1}_{(\theta,\infty)}(x_{(1)})\,\,(i)$
is a joint density of $(X_{1},\cdots,X_{n})$ w.r.t. the Lebesgue measure on $\mathbb{R}^{n}$. Using the Neyman-Fisher Factorization Theorem, we infer that $X_{(1)}$ is sufficient. Define $\Theta_{0}:=\mathbb{Q}$. Then $\Theta_{0}\subseteq\mathbb{R}$ is countable. For each $x\in\mathbb{R}^{n}$, define
$D(x,\Theta_{0}):=\big\{y\in\mathbb{R}^{n}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta_{0})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}.$
Let $x,y\in\mathbb{R}^{n}$ be arbitrary elements such that $y\in D(x,\Theta_{0})\,\,(ii)$. Because of Method 3.1, to prove that $X_{(1)}$ is minimal sufficient, it suffices to prove that $x_{(1)}=y_{(1)}$. From $(ii)$, we know that there exists a finite constant $h_{xy}>0$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}\,\,(iii)$ for every $\theta\in\Theta_{0}$. Suppose, by contradiction, that $x_{(1)}\neq y_{(1)}$. Then $x_{(1)}<y_{(1)}$ or $y_{(1)}<x_{(1)}$. Suppose that $x_{(1)}<y_{(1)}$. Then there exists a rational number $\theta_{0}\in\Theta_{0}:=\mathbb{Q}$ such that $x_{(1)}<\theta_{0}<y_{(1)}\,\,(iv)$. From $(iii)$, we know that $f_{\theta_{0}}(y)=f_{\theta_{0}}(x)h_{xy}$, which implies, using $(i)$ and $(iv)$, that $f_{\theta_{0}}(y)=f_{\theta_{0}}(x)h_{xy}=0$, since $f_{\theta_{0}}(x)=0$. However, this is a contradiction since $f_{\theta_{0}}(y)>0$. Therefore, the assumption that $x_{(1)}<y_{(1)}$ is false. Analogously, we can conclude that $y_{(1)}<x_{(1)}$ does not hold, implying that $x_{(1)}=y_{(1)}$.
Example $3.$
Let $X_{1},\cdots,X_{n}$ be a random sample with density $p_{\theta}$ given by
$p_{\theta}(x)=\left(\frac{2}{\pi}\right)^{1/2}\exp\left\{-\frac{(x-\theta)^{2}}{2}\right\}\mathbf{1}_{[\theta,\infty)}(x),\,\,\,\,\theta\in\mathbb{R}.$
We will prove that $(\overline{X},X_{(1)})$ is a minimal sufficient statistic. The function $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ given by $f_{\theta}(x_{1},\cdots,x_{n})=(2/\pi)^{n/2}e^{n\theta\overline{x}-n\theta^{2}/2}e^{-(\sum_{i=1}^{n}x_{i}^{2})/2}\mathbf{1}_{[\theta,\infty)}(x_{(1)})\,\,(i)$ is a joint density of $(X_{1},\cdots,X_{n})$ w.r.t. the Lebesgue measure on $\mathbb{R}^{n}$. Using the Neyman-Fisher Factorization Theorem, we conclude that $(\overline{X},X_{(1)})$ is a sufficient statistic. Define $\Theta_{0}:=\mathbb{Q}$. Then $\Theta_{0}\subseteq\mathbb{R}$ is countable. For each $x\in\mathbb{R}^{n}$, define
$D(x,\Theta_{0}):=\big\{y\in\mathbb{R}^{n}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta_{0})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}.$
Let $x,y\in\mathbb{R}^{n}$ be arbitrary elements such that $y\in D(x,\Theta_{0})\,\,(ii)$. Because of Method 3.1, to prove that $(\overline{X},X_{(1)})$ is minimal sufficient, it suffices to prove that $(\overline{x},x_{(1)})=(\overline{y},y_{(1)})$. From $(ii)$, we know that there exists a finite constant $h_{xy}>0$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\Theta_{0}$. Therefore, using $(i)$, we conclude that for every $\theta\in\Theta_{0}$, we have
$e^{n\theta\overline{y}}e^{-(\sum_{i=1}^{n}y_{i}^{2})/2}\mathbf{1}_{[\theta,\infty)}(y_{(1)})=e^{n\theta\overline{x}}e^{-(\sum_{i=1}^{n}x_{i}^{2})/2}\mathbf{1}_{[\theta,\infty)}(x_{(1)})h_{xy}\,\,(iii).$
Repeating an argument from Example 3.2, we can conclude that $x_{(1)}=y_{(1)}$, which implies, using $(iii)$, that, for each $\theta\in\Theta_{0}\cap(-\infty,x_{(1)}]$, we have $e^{n\theta(\overline{y}-\overline{x})}=e^{-(\sum_{i=1}^{n}x_{i}^{2})/2}h_{xy}e^{(\sum_{i=1}^{n}y_{i}^{2})/2}$. Since the previous equality holds for every $\theta\in\Theta_{0}\cap(-\infty,x_{(1)}]$ and its right side does not depend on $\theta$, we can conclude that $\overline{y}-\overline{x}=0$ and, consequently, that $\overline{x}=\overline{y}$. Hence, $(\overline{x},x_{(1)})=(\overline{y},y_{(1)})$.
Example 3.4.
Let $(X_{1},Y_{1}),\cdots,(X_{n},Y_{n})$ be a random sample with density $p_{\theta}(x,y)=2/\theta^{2}$, for $x>0$, $y>0$, and $x+y<\theta$, and $p_{\theta}(x,y)=0$, otherwise, with $\theta>0$. Define $X:=(X_{1},\cdots,X_{n})$ and $Y:=(Y_{1},\cdots,Y_{n})$. We will prove that the order statistic $(X+Y)_{(n)}$ is minimal sufficient. The function $f_{\theta}:((0,\infty)^{2})^{n}\to\mathbb{R}$ given by
$f_{\theta}((x_{1},y_{1}),\cdots,(x_{n},y_{n})):=2^{n}\theta^{-2n}\mathbf{1}_{(0,\infty)^{2}}(x_{(1)},y_{(1)})\mathbf{1}_{(-\infty,\theta)}((x+y)_{(n)})\,\,(i)$
is a joint density of $((X_{1},Y_{1}),\cdots,(X_{n},Y_{n}))$ w.r.t. the Lebesgue measure on $((0,\infty)^{2})^{n}$. Using Propositions 8.1.2, 8.1.3, and 8.1.7 of Cohn (2013), we can conclude that the product $((0,\infty)^{2})^{n}$ of $n$ copies of $(0,\infty)^{2}$ together with its Borel $\sigma$-algebra is a standard Borel space. Using $(i)$ and the Neyman-Fisher Factorization Theorem, we can conclude that the statistic $(X+Y)_{(n)}$ is sufficient. Define $\Theta_{0}:=\mathbb{Q}\cap(0,\infty)$. Then $\Theta_{0}\subseteq(0,\infty)$ is countable. For each $a\in((0,\infty)^{2})^{n}$, define $D(a,\Theta_{0}):=\big\{b\in((0,\infty)^{2})^{n}:(\exists h_{ab}\in(0,\infty))(\forall\theta\in\Theta_{0})\big(f_{\theta}(b)=f_{\theta}(a)h_{ab}\big)\big\}$. Let $a:=((x_{1},y_{1}),\cdots,(x_{n},y_{n})),b:=((z_{1},w_{1}),\cdots,(z_{n},w_{n}))\in((0,\infty)^{2})^{n}$ be arbitrary such that $b\in D(a,\Theta_{0})\,\,(ii)$. Because of Method 3.1, to prove that $(X+Y)_{(n)}$ is minimal sufficient, it suffices to prove that $(x+y)_{(n)}=(z+w)_{(n)}$, where $x:=(x_{1},\cdots,x_{n})$ and $y,z,w$ are defined likewise. From $(ii)$, we know that there exists a finite constant $h_{ab}>0$ such that $f_{\theta}(b)=f_{\theta}(a)h_{ab}$ for every $\theta\in\Theta_{0}$. Then, using $(i)$, we can conclude that, for each $\theta\in\Theta_{0}$, we have
$\mathbf{1}_{(-\infty,\theta)}((z+w)_{(n)})=\mathbf{1}_{(-\infty,\theta)}((x+y)_{(n)})h_{ab}\,\,(iii).$
Suppose, by contradiction, that $(x+y)_{(n)}<(w+z)_{(n)}$. Then, using $(iii)$, we obtain $0=h_{ab}$ for every $\theta\in((x+y)_{(n)},(z+w)_{(n)})\cap\Theta_{0}$, a contradiction. Hence, $(w+z)_{(n)}\leq(x+y)_{(n)}$. Analogously, we can conclude that $(x+y)_{(n)}\leq(w+z)_{(n)}$, implying that $(x+y)_{(n)}=(z+w)_{(n)}$.
Recall that, in Method 3.1, the implication $y\in D(x,\Theta_{0})\ \Longrightarrow\ T(x)=T(y)$ must hold for all $x,y\in\mathcal{X}$. This makes the method not directly applicable to some statistics, as the following example shows.
Example 3.5.
Let $(X_{1},X_{2})$ be a vector with density $f_{\theta}(x_{1},x_{2})=(4/\pi)\theta^{3}x_{1}^{2}x_{2}^{2}\exp(-\theta(x_{1}^{2}+x_{2}^{2}))$, for $x_{1},x_{2}\in\mathbb{R}$, and $\theta>0$. By the Neyman-Fisher Factorization Theorem, $T(X_{1},X_{2})=X_{1}^{2}+X_{2}^{2}$ is sufficient. Note that $f_{\theta}(x_{1},x_{2})=0$ whenever $x_{1}x_{2}=0$. Take, for instance, $x=(0,2)$ and $y=(0,1)$. Then $T(x)=4\neq T(y)=1$, but, as $f_{\theta}(x)=f_{\theta}(y)=0$ for every $\theta>0$, we have, for any countable set $\Theta_{0}\subset(0,\infty)$, that $y\in D(x,\Theta_{0})$. Therefore, Method 3.1 cannot be applied directly on $T$, since the required implication $y\in D(x,\Theta_{0})$ implies T(x)=T(y)T(x)=T(y)” fails. However, we can define a pointwise modified statistic which is equal to $T$ $P_{\theta}$-a.e. for each $\theta>0$. For example, let $\tilde{T}(x_{1},x_{2})=T(x_{1},x_{2})$, if $x_{1}x_{2}\neq 0$, and $\tilde{T}(x_{1},x_{2})=1$, otherwise. Now let $\Theta_{0}:=\mathbb{Q}\cap(0,\infty)$ and assume that $y\in D(x,\Theta_{0})$. Then there exists $h_{xy}\in(0,\infty)$ such that $f_{\theta}(y)=f_{\theta}(x)\,h_{xy}$, $\forall\theta\in\Theta_{0}$.
There are two cases, namely, (1) if $x_{1}x_{2}=0$, then $f_{\theta}(x)=0$ for every $\theta\in\Theta_{0}$. Hence $f_{\theta}(y)=0$ for every $\theta\in\Theta_{0}$ as well, so necessarily $y_{1}y_{2}=0$. Therefore, $\tilde{T}(x)=1=\tilde{T}(y)$; and (2) if $x_{1}x_{2}\neq 0$, then $f_{\theta}(x)>0$ for every $\theta\in\Theta_{0}$. Since $y\in D(x,\Theta_{0})$ and $h_{xy}>0$, we must also have $y_{1}y_{2}\neq 0$. In this case,
$h_{xy}=\frac{f_{\theta}(y)}{f_{\theta}(x)}=\frac{y_{1}^{2}y_{2}^{2}}{x_{1}^{2}x_{2}^{2}}\exp\big(-\theta(\tilde{T}(y)-\tilde{T}(x))\big),\qquad\forall\theta\in\Theta_{0}.$
Because $h_{xy}$ does not depend on $\theta$ and $\Theta_{0}$ contains infinitely many distinct values, we must have $\tilde{T}(x)=\tilde{T}(y)$. Thus the implication required in Method 3.1 holds for $\tilde{T}$. Since $\tilde{T}=T$ $P_{\theta}$-a.e. for every $\theta>0$, $\tilde{T}$ is also sufficient. Hence, by Method 3.1, $\tilde{T}$ is minimal sufficient. Moreover, since $T=\tilde{T}$ $P_{\theta}$-almost surely for every $\theta>0$, $T$ is also minimal sufficient.
The following method is a generalization of the one introduced by Sato (1996), whose original proof was restricted to Euclidean spaces. Its formal proof is deferred to Section 4.
Method 3.2 (Sato).
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space, $(\mathcal{T},\Sigma_{\mathcal{T}})$ is a separable measurable space, and $T:\mathcal{X}\to\mathcal{T}$ is a measurable function. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Suppose that there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that, for each $\theta\in\Theta$, there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ such that the limit $\lim_{n\to\infty}f_{\theta_{n}}$ exists in $\mathbb{R}$ $\mu$-a.e. and for every $x\in\mathcal{X}$ we have
$f_{\theta}(x)=\begin{cases}\lim_{n\to\infty}f_{\theta_{n}}(x),&\text{if the limit exists}\\ 0,&\text{otherwise}\end{cases}.$
Define $D(x):=\big\{y\in\mathcal{X}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta)\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}$. If for any $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff $y\in D(x)$, then $T$ is minimal sufficient.
This method states that, under the approximation condition above, the usual likelihood-ratio characterization of minimal sufficiency becomes valid: $T$ is minimal sufficient if, for any sample points $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff there exists a finite constant $h_{xy}>0$ independent of $\theta$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\Theta$.
Note that the latter approach highlights the necessary conditions under which Criterion 1.1 introduced in Section 1 becomes valid. It is particularly useful when $\Theta$ is a subset of $\mathbb{R}^{n}$ and the densities $f_{\theta}$ are continuous w.r.t. $\theta$, but it is difficult to apply in cases such as Example 3.1. The following example illustrates its application.
Example 3.6.
Let $X_{1},\cdots,X_{n}\sim\text{Cauchy}(\theta,1),\theta\in\mathbb{R}$ be a random sample. The function $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ given by $f_{\theta}(x_{1},\cdots,x_{n}):=\pi^{-n}\prod_{i=1}^{n}\frac{1}{1+(x_{i}-\theta)^{2}}$ is a joint density of $(X_{1},\cdots,X_{n})$ w.r.t. the Lebesgue measure on $\mathbb{R}^{n}$. We will show that $(X_{(1)},X_{(2)},\cdots,X_{(n)})$ is a minimal sufficient statistic, where $X_{(k)}$ denotes the $k$-th order statistic. Since the function $\mathbb{R}\to\mathbb{R},\,\theta\mapsto f_{\theta}(x)$ is continuous for every $x\in\mathbb{R}^{n}$ and the set of rational numbers $\mathbb{Q}$ is countable and dense in $\mathbb{R}$, the conditions of Method 3.2 are satisfied. Define
$D(x):=\big\{y\in\mathbb{R}^{n}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\mathbb{R})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}$
for every $x\in\mathbb{R}^{n}$. Since $h/f_{\theta}(x)$ is a polynomial in $\theta$ for every $h>0$ and $x:=(x_{1},\cdots,x_{n})\in\mathbb{R}^{n}$ with complex roots $x_{j}\pm i,j\in\{1,\cdots,n\}$, as two identical polynomials have the same roots, it is straightforward to show that, for any two sample points $x,y\in\mathbb{R}^{n}$, we have $(x_{(1)},\cdots,x_{(n)})=(y_{(1)},\cdots,y_{(n)})$ iff $y\in D(x)$. Hence, using Method 3.2, we infer that $(X_{(1)},X_{(2)},\cdots,X_{(n)})$ is minimal sufficient.
We conclude this section with a method for exponential models, based on Proposition 1.6.9 in Pfanzagl (1994). Pfanzagl’s original proof invokes his minimality criterion (cf. Criterion 1.2), which we showed to be false without extra assumptions in Section 2. We provide a complete proof of the following method in Section 4; our version holds under slightly stronger hypotheses than Pfanzagl’s original formulation.
Method $3.$
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $P_{\theta}\ll\mu$ for every $\theta\in\Theta$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space. For each $\theta\in\Theta$, let $f_{\theta}:\mathcal{X}\to\mathbb{R}$ be a density of $P_{\theta}$ w.r.t. $\mu$. Suppose that there exist functions $\eta_{1},\cdots,\eta_{k},B:\Theta\to\mathbb{R}$ and measurable functions $T_{1},\cdots,T_{k},h:\mathcal{X}\to\mathbb{R}$ such that, for each $\theta\in\Theta$, we have
$f_{\theta}(x)=\exp\big[\big(\sum_{i=1}^{k}\eta_{i}(\theta)T_{i}(x)\big)-B(\theta)\big]h(x)\quad\mu\text{-a.e.}$
If for any $a_{0},\cdots,a_{k}\in\mathbb{R}$ satisfying $\sum_{i=1}^{k}a_{i}\eta_{i}(\theta)=a_{0}$ for every $\theta\in\Theta$, we have $a_{i}=0$ for every $i\in\{0,\cdots,k\}$, then $T:\mathcal{X}\to\mathbb{R}^{k}$ given by $T(x):=(T_{1}(x),\cdots,T_{k}(x))$ is a minimal sufficient statistic.
The following examples illustrate the application of this method.
Example 3.7.
Let $X_{1},\cdots,X_{n}\sim N(\theta,k\theta^{2}),\theta>0$ be a random sample, where $k>0$ is a known constant. We will prove that $(\sum_{i=1}^{n}X_{i},\sum_{i=1}^{n}X_{i}^{2})$ is a minimal sufficient statistic. The function $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ given by
$f_{\theta}(x_{1},\cdots,x_{n}):=\frac{e^{-n/(2k)}}{(2\pi k)^{n/2}}\exp\left\{\frac{1}{k\theta}\left(\sum_{i=1}^{n}x_{i}\right)-\frac{1}{2k\theta^{2}}\left(\sum_{i=1}^{n}x_{i}^{2}\right)-n\log(\theta)\right\}$
is a joint density of $(X_{1},\cdots,X_{n})$ w.r.t. the Lebesgue measure on $\mathbb{R}^{n}$.
For each $\theta>0$, define $\eta_{1}(\theta):=1/(k\theta)$ and $\eta_{2}(\theta):=-1/(2k\theta^{2})$. Let $a,b,c\in\mathbb{R}$ be arbitrary elements such that $a\eta_{1}(\theta)+b\eta_{2}(\theta)=c\,\,(i)$ for every $\theta>0$. Because of Method 3.3, to prove that $(\sum_{i=1}^{n}X_{i},\sum_{i=1}^{n}X_{i}^{2})$ is minimal sufficient, it suffices to prove that $a=b=c=0$. From $(i)$, we know that $2kc\theta^{2}-2a\theta+b=0$ for every $\theta>0$, allowing us to conclude, noting that this is a polynomial with degree at most 22 in $\theta$, that 2kc=2a=b=02kc=-2a=b=0 and, hence, that $a=b=c=0$.
4 Proofs
Before proving the main results, it is necessary to establish some preliminaries.
Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model. Whenever we write $f:\mathcal{X}\to(\mathcal{Y},\Sigma_{\mathcal{Y}})$, we mean that $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ is a measurable space and that $f:\mathcal{X}\to\mathcal{Y}$ is $(\Sigma,\Sigma_{\mathcal{Y}})$-measurable. Moreover, when $f:\mathcal{X}\to\mathcal{Y}$ is $(\mathfrak{B},\Sigma_{\mathcal{Y}})$-measurable, where $\mathfrak{B}$ is a sub-sigma-algebra of $\Sigma$, we say that $f$ is $\mathfrak{B}$-measurable.
Let $\mathcal{X}$ be a set and $\{\mathfrak{A}_{i}\}_{i\in I}$ a collection of subsets of $2^{\mathcal{X}}$. We denote by $\vee_{i\in I}\mathfrak{A}_{i}$ the $\sigma$-algebra on $\mathcal{X}$ generated by the union $\cup_{i\in I}\mathfrak{A}_{i}$. Furthermore, we denote $\vee_{i\in I}\mathfrak{A}_{i}$ by $\mathfrak{A}_{1}\vee\cdots\vee\mathfrak{A}_{n}$ if $I=\{1,\cdots,n\}$.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mathfrak{A}$ a sub-$\sigma$-algebra of $\Sigma$. We denote
$\mathfrak{N}_{\mathcal{E}}:=\big\{N\in\Sigma:(\forall\theta\in\Theta)\big(P_{\theta}(N)=0\big)\big\}.$
That is, $\mathfrak{N}_{\mathcal{E}}$ is the set of all $N\in\Sigma$ that are null with respect to every probability measure in the statistical model $\mathcal{E}$. Moreover, we denote $\overline{\mathfrak{A}}^{\mathcal{E}}:=\mathfrak{A}\vee\mathfrak{N}_{\mathcal{E}}$ for every sub-$\sigma$-algebra $\mathfrak{A}\subseteq\Sigma$. Hence, $\overline{\mathfrak{A}}^{\mathcal{E}}$ is the $\sigma$-algebra on $\mathcal{X}$ generated by the union $\mathfrak{A}\cup\mathfrak{N}_{\mathcal{E}}$. Repeating the argument in Lemma 4.1.3 of Doberkat (2015), we obtain
$\overline{\mathfrak{A}}^{\mathcal{E}}=\big\{A\vartriangle N:A\in\mathfrak{A}\wedge N\in\mathfrak{N}_{\mathcal{E}}\big\}.$
Therefore, given two sub-$\sigma$-algebras $\mathfrak{A}$ and $\mathfrak{B}$ of $\Sigma$, we have $\mathfrak{A}\subseteq\overline{\mathfrak{B}}^{\mathcal{E}}$ iff for each $A\in\mathfrak{A}$ there exists $B\in\mathfrak{B}$ such that $P_{\theta}(A\vartriangle B)=0$ for every $\theta\in\Theta$.
Proposition 4.1.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model, $\mathfrak{A}$ a sub-$\sigma$-algebra of $\Sigma$, and $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ a standard Borel space. If $f:\mathcal{X}\to\mathcal{Y}$ is $\overline{\mathfrak{A}}^{\mathcal{E}}$-measurable, then there exists a $\mathfrak{A}$-measurable function $f_{0}:\mathcal{X}\to\mathcal{Y}$ such that, for every $\theta\in\Theta$, we have $f=f_{0}$ $P_{\theta}$-a.e.
Proof.
See Lemma 1.10.3, p. 56, in Pfanzagl (1994)
The previous proposition can be generalized to separable measurable spaces, as can be seen in Cavalcante (2026, Proposition 2.1.4).
Let $\mathcal{M}_{1}$, $\mathcal{M}_{2}$ and $\mathcal{M}$ be collections of measures on a set $(\mathcal{X},\Sigma)$. We denote $\mathcal{M}_{1}\ll\mathcal{M}_{2}$ iff for every $E\in\Sigma$ satisfying $(\forall\nu\in\mathcal{M}_{2})(\nu(E)=0)$ we have $(\forall\mu\in\mathcal{M}_{1})(\mu(E)=0)$. We also denote $\mathcal{M}_{1}\equiv\mathcal{M}_{2}$ iff $\mathcal{M}_{1}\ll\mathcal{M}_{2}$ and $\mathcal{M}_{2}\ll\mathcal{M}_{1}$. Moreover, we denote $\mathcal{E}\ll\mathcal{M}$ (resp. $\mathcal{E}\equiv\mathcal{M}$) iff $\{P_{\theta}\}_{\theta\in\Theta}\ll\mathcal{M}$ (resp. $\{P_{\theta}\}_{\theta\in\Theta}\equiv\mathcal{M}$), where $\mathcal{E}$ is the statistical model. If $\mathcal{M}=\{\mu\}$ is a singleton, we denote $\mathcal{E}\ll\mathcal{M}$ and $\mathcal{E}\equiv\mathcal{M}$ simply by $\mathcal{E}\ll\mu$ and $\mathcal{E}\equiv\mu$, respectively.
Proposition 4.2.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu\!:\!\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure. If $\mathcal{E}\ll\mu$, then there exists a countable subset $\{\theta_{n}\}_{n\in\mathbb{N}}\subseteq\Theta$ such that for every sequence $(c_{n})_{n\in\mathbb{N}}$ of $(0,\infty)$ satisfying $\sum_{n\in\mathbb{N}}c_{n}=1$, we have that $P:\Sigma\to\mathbb{R}$ given by ${P}:=\sum_{n\in\mathbb{N}}c_{n}P_{\theta_{n}}$ is a probability measure such that $\mathcal{E}\equiv{P}$, $P\ll\mu$, and
$\frac{d{P}}{d\mu}=\sum_{n\in\mathbb{N}}c_{n}\frac{dP_{\theta_{n}}}{d\mu}\,\,\mu\text{-a.e.}$
Proof.
Suppose that $\mathcal{E}\ll\mu$. From Lemma 7 in Halmos and Savage (1949), we know that there exists a countable subset $\{\theta_{n}\}_{n\in\mathbb{N}}\subseteq\Theta$ such that $\mathcal{E}\equiv\{P_{\theta_{n}}\}_{n\in\mathbb{N}}$. Let $(c_{n})_{n\in\mathbb{N}}$ be an arbitrary sequence of $(0,\infty)$ such that $\sum_{n\in\mathbb{N}}c_{n}=1$. Define $P:\Sigma\to\mathbb{R}$ by ${P}:=\sum_{n\in\mathbb{N}}c_{n}P_{\theta_{n}}$. Then $P$ is a probability measure. Moreover, we have ${P}\ll\{P_{\theta_{n}}\}_{n\in\mathbb{N}}\ll\{P_{\theta}\}_{\theta\in\Theta}\ll\{P_{\theta_{n}}\}_{n\in\mathbb{N}}\ll{P}$, implying that $\mathcal{E}\equiv{P}$.
By hypothesis, $\mathcal{E}\ll\mu$; therefore, $\{P_{\theta_{n}}\}_{n\in\mathbb{N}}\ll\mu$ and ${P}\ll\mu$. Hence, given the almost everywhere uniqueness of a Radon-Nikodym derivative, it is direct to verify that $\frac{d{P}}{d\mu}=\sum_{n\in\mathbb{N}}c_{n}\frac{dP_{\theta_{n}}}{d\mu}$ $\mu$-a.e. ∎
Let $\mu,\nu$ be two measures on $(\mathcal{X},\Sigma)$ such that $\nu\ll\mu$. We denote by $d\nu/d\mu]$ the set of all Radon-Nikodym derivatives of $\nu$ w.r.t. $\mu$. Define $d_{\mathcal{E}}:\Theta\times\Theta\to\mathbb{R}$, the total variation distance, by $d_{\mathcal{E}}(\theta_{1},\theta_{2}):=\sup_{E\in\Sigma}|P_{\theta_{1}}(E)-P_{\theta_{2}}(E)|.$
Proposition 4.3.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model, $\theta_{1},\theta_{2}\in\Theta$, and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $\mathcal{E}\ll\mu$. If $f_{\theta_{1}}\in[dP_{\theta_{1}}/d\mu]$ and $f_{\theta_{2}}\in[dP_{\theta_{2}}/d\mu]$, then
$d_{\mathcal{E}}(\theta_{1},\theta_{2})=\frac{1}{2}\int_{\mathcal{X}}|f_{\theta_{1}}-f_{\theta_{2}}|d\mu.$
Proof.
See Lemma 2.4 in Strasser (1985). ∎
We denote by $\tau\langle\mathcal{E}\rangle$ the coarsest topology on $\Theta$ that makes the functions $\Phi_{E}:\Theta\to\mathbb{R}$ given by $\Phi_{E}(\theta):=P_{\theta}(E)$ continuous for every $E\in\Sigma$. Hence, $\tau\langle\mathcal{E}\rangle$ is the initial topology on $\Theta$ induced by the collection $\{\Phi_{E}:\Theta\to\mathbb{R}\}_{E\in\Sigma}$.
Let $(\mathcal{X},d)$ be a pseudo-metric space. We denote by $\tau(d)$ the topology on $\mathcal{X}$ whose elements are arbitrary unions of open balls defined by $d$. That is, $\tau(d)$ is the topology on $\mathcal{X}$ induced by the pseudo-metric $d$.
Proposition $4.$
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model. Then
$\tau\langle\mathcal{E}\rangle\subseteq\tau(d_{\mathcal{E}}).$
Proof.
For each $E\in\Sigma$, define $\Phi_{E}:\Theta\to\mathbb{R}$ by $\Phi_{E}(\theta):=P_{\theta}(E)$. To conclude the proof, it suffices to prove that $\Phi_{E}$ is $\tau(d_{\mathcal{E}})$-continuous. This follows from the observation that for any $E\in\Sigma$ and $\theta,\theta_{0}\in\Theta$ we have $|\Phi_{E}(\theta)-\Phi_{E}(\theta_{0})|=|P_{\theta}(E)-P_{\theta_{0}}(E)|\leq d_{\mathcal{E}}(\theta,\theta_{0})$, by definition of the total variation distance. ∎
An immediate consequence of the previous proposition is the next corollary.
Corollary 4.1.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\Theta_{0}\subseteq\Theta$. If $\Theta_{0}$ is dense in $(\Theta,d_{\mathcal{E}})$, then $\Theta_{0}$ is dense in $(\Theta,\tau\langle\mathcal{E}\rangle)$.
Before stating the next proposition, we extend the notion of (minimal) sufficiency to $\sigma$-algebras. Let $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model. A sub-$\sigma$-algebra $\mathfrak{A}$ of $\Sigma$ is said to be ($\mathcal{E}$-)sufficient (or $\{P_{\theta}\}_{\theta\in\Theta}$-sufficient) iff for each $E\in\Sigma$, there exists a $\mathfrak{A}$-measurable function $\kappa_{E}:\mathcal{X}\to\mathbb{R}$ such that for every $\theta\in\Theta$ we have $P\!_{\theta}(E|\mathfrak{A})=\kappa_{E}$ $P_{\theta}$-a.e. Furthermore, $\mathfrak{A}$ is said to be minimal ($\mathcal{E}$-)sufficient (or minimal $\{P_{\theta}\}_{\theta\in\Theta}$-sufficient) iff $\mathfrak{A}$ is $\mathcal{E}$-sufficient and, given any $\mathcal{E}$-sufficient sub-$\sigma$-algebra $\mathfrak{B}$ of $\Sigma$, we have $\mathfrak{A}\subseteq\overline{\mathfrak{B}}^{\mathcal{E}}$.
There exists a connection between sufficient statistics and sufficient $\sigma$-algebras. Before explicating this relation, we require the following lemma.
Proposition 4.5 (Doob-Dynkin Lemma).
Let $\mathcal{X}$ be a set, $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ and $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ two measurable spaces, and $f:\mathcal{X}\to\mathcal{Y}$ and $g:\mathcal{X}\to\mathcal{Z}$ two functions. If $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ is a standard Borel space, then the following statements are equivalent:
  1. 1.
    $\sigma(g)\subseteq\sigma(f)$;
  2. 2.
    There exists a measurable function $h:\mathcal{Y}\to\mathcal{Z}$ such that $g=h\circ f$.
Proof.
See Lemma 1.14, p. 18, in Kallenberg (2021)
Using the Doob-Dynkin Lemma, it is straightforward to show that a statistic $T:\mathcal{X}\to(\mathcal{T},\Sigma_{\mathcal{T}})$ is $\mathcal{E}$-sufficient iff $\sigma(T)$ is $\mathcal{E}$-sufficient. This equivalence generally does not extend to minimal sufficiency; as demonstrated in Examples 9.7 and 9.8 of Heyer (1982), $T$ may be minimal sufficient while $\sigma(T)$ is not, and conversely. The following proposition establishes a sufficient assumption on the statistic’s codomain under which this discrepancy disappears in one direction: if the codomain is standard Borel (see Definition A.3), then minimality at the $\sigma$-algebra level lifts to minimality at the statistic level.
Proposition 4.6.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $T:\mathcal{X}\to(\mathcal{T},\Sigma_{\mathcal{T}})$. Suppose that $(\mathcal{T},\Sigma_{\mathcal{T}})$ is a standard Borel space. If $\sigma(T)$ is minimal sufficient, then $T$ is minimal sufficient.
Proof.
Suppose that $\sigma(T)$ is minimal sufficient. Then $\sigma(T)$ is sufficient, implying that $T$ is sufficient. Let $S:\mathcal{X}\to(\mathcal{S},\Sigma_{\mathcal{S}})$ be an arbitrary sufficient statistic. Then $\sigma(S)$ is sufficient, which implies, using the hypothesis, that $\sigma(T)\subseteq\overline{\sigma(S)}^{\mathcal{E}}$. Therefore, $T$ is $\overline{\sigma(S)}^{\mathcal{E}}$-measurable, which implies, using Proposition 4.1, that there exists a $\sigma(S)$-measurable function $g:\mathcal{X}\to\mathcal{T}$ such that, for each $\theta\in\Theta$, we have $T=g$ $P_{\theta}$-a.e.
Since $g$ is $\sigma(S)$-measurable, we have $\sigma(g)\subseteq\sigma(S)$, which implies, using the Doob-Dynkin Lemma (see Proposition 4.5), that there exists a $(\Sigma_{\mathcal{S}},\Sigma_{\mathcal{T}})$-measurable function $f:\mathcal{S}\to\mathcal{T}$ such that $g=f(S)$, implying that, for each $\theta\in\Theta$, we have $T=f(S)$ $P_{\theta}$-a.e. ∎
Proposition 4.7.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $(c_{n})_{n\in\mathbb{N}}$ a sequence of $(0,\infty)$ such that $\sum_{n\in\mathbb{N}}c_{n}=1$. If $\{\theta_{n}\}_{n\in\mathbb{N}}\subseteq\Theta$ is dense in $(\Theta,\tau\langle\mathcal{E}\rangle)$, then the following statements are true:
  1. i.
    $P:\Sigma\to\mathbb{R}$ given by $P:=\sum_{n\in\mathbb{N}}c_{n}P_{\theta_{n}}$ is a probability measure such that $\mathcal{E}\equiv P$;
  2. ii.
    For each $n\in\mathbb{N}$, choose $f_{\theta_{n}}\in[dP_{\theta_{n}}/dP]$. Then $\mathfrak{A}:=\vee_{n\in\mathbb{N}}\sigma(f_{\theta_{n}})$ is a countably generated minimal $\mathcal{E}$-sufficient sub-$\sigma$-algebra of $\Sigma$.
Proof.
Suppose that $\{\theta_{n}\}_{n\in\mathbb{N}}\subseteq\Theta$ is dense in $(\Theta,\tau\langle\mathcal{E}\rangle)$. Define $P:\Sigma\to\mathbb{R}$ by $P:=\sum_{n\in\mathbb{N}}c_{n}P_{\theta_{n}}$. Then $P$ is a probability measure. Repeating an argument from Lemma 4.3 of Strasser (1985), we conclude that $\mathcal{E}\ll P$, which implies that $\mathcal{E}\equiv P$ since $P\ll\{P_{\theta_{n}}\}_{n\in\mathbb{N}}$. Hence, $\mathcal{E}$ is dominated.
For each $\theta\in\Theta$, choose $f_{\theta}\in[dP_{\theta}/dP]$. Define $\mathfrak{A}:=\vee_{n\in\mathbb{N}}\sigma(f_{\theta_{n}})$.
Repeating an argument from that lemma, we conclude that $\mathfrak{A}$ is a countably generated sub-$\sigma$-algebra of $\Sigma$ and that $f_{\theta}=\mathbb{E}_{P}[f_{\theta}|\mathfrak{A}]$ $P$-a.e. for every $\theta\in\Theta$, allowing us to conclude that $\mathfrak{A}$ is sufficient (Strasser, 1985, Lemma 20.6). Hence, it remains to prove that $\mathfrak{A}$ is minimal.
From the proof of Theorem 8.8 in Heyer (1982), we know $\mathfrak{B}:=\vee_{\theta\in\Theta}\sigma(f_{\theta})$ is minimal sufficient, which allows us to conclude that $\mathfrak{A}$ is minimal since $\mathfrak{A}\subseteq\mathfrak{B}$.
A more detailed proof can be found in Proposition 2.5.5 of Cavalcante (2026). ∎
We now state a result similar to the Doob-Dynkin Lemma that will be employed to prove our proposed methods.
Proposition 4.8.
Let $(\mathcal{X},\Sigma_{\mathcal{X}})$, $(\mathcal{Y},\Sigma_{\mathcal{Y}})$, and $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ be three measurable spaces and $f:\mathcal{X}\to\mathcal{Y}$ and $g:\mathcal{X}\to\mathcal{Z}$ two measurable functions. Suppose that $(\mathcal{X},\Sigma_{\mathcal{X}})$ is an analytic Borel space and $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ is separable. If $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$, then $\sigma(g)\subseteq\sigma(f)$.
Proof.
Suppose that $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$. Then $h:f[\mathcal{X}]\to\mathcal{Z}$ given by $h(f(x)):=g(x)$ is well-defined.
Let $E\in\sigma(g)$ be an arbitrary element. Then there exists $F\in\Sigma_{\mathcal{Z}}$ such that $E=g^{-1}[F]$, allowing us to conclude that
$E=(h\circ f)^{-1}[F]=f^{-1}[h^{-1}[F]]=\cup_{y\in h^{-1}[F]}f^{-1}[\{y\}].$ (1)
By hypothesis, $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ is countably generated; therefore, $\sigma(f)$ is countably generated (since if $\Sigma_{\mathcal{Y}}$ is generated by $\mathcal{S}\subseteq 2^{\mathcal{Y}}$, then $\sigma(f)$ is generated by $f^{-1}(\mathcal{S})$). Moreover, it is straightforward to verify that $f^{-1}[\{y\}]$ is an atom of $\sigma(f):=f^{-1}(\Sigma_{\mathcal{Y}})$ for every $y\in f[\mathcal{X}]$ since, by hypothesis, $\Sigma_{\mathcal{Y}}$ is separable (see Section 8.6 of Cohn, 2013, for the definition of an atom).
As $f$ is measurable, it follows that $\sigma(f)$ is a sub-$\sigma$-algebra of $\Sigma_{\mathcal{X}}$ and $E\in\Sigma_{\mathcal{X}}$. Recall that if $(\mathcal{W},\Sigma)$ is a separable measurable space, then $(A,\Sigma|_{A})$ is separable for every subset $A\subseteq\mathcal{W}$. Hence, by Proposition 8.6.5 of Cohn (2013), our definition of an analytic Borel space coincides with the definition of an analytic space given in Section 8.6 of Cohn (2013). Therefore, using (1) and Theorem 8.6.7 of Cohn (2013), we conclude $E\in\sigma(f)$. ∎
Proposition 4.9.
Let $(\mathcal{X},\Sigma_{\mathcal{X}})$, $(\mathcal{Y},\Sigma_{\mathcal{Y}})$, and $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ be three measurable spaces and $f:\mathcal{X}\to\mathcal{Y}$ and $g:\mathcal{X}\to\mathcal{Z}$ two measurable functions. If $(\mathcal{X},\Sigma_{\mathcal{X}})$ is an analytic Borel space, $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ is separable, and $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ is a standard Borel space, then the following statements are equivalent:
  1. 1.
    $\sigma(g)\subseteq\sigma(f)$;
  2. 2.
    There exists a measurable function $h:\mathcal{Y}\to\mathcal{Z}$ such that $g=h\circ f$;
  3. 3.
    There exists a (not necessarily measurable) function $\tilde{h}:\mathcal{Y}\to\mathcal{Z}$ such that $g=\tilde{h}\circ f$;
  4. 4.
    $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$.
Proof.
The equivalence $1.\Leftrightarrow 2.$ is a direct consequence of the Doob-Dynkin Lemma. The implications $2.\Rightarrow 3.$ and $3.\Rightarrow 4.$ are trivial. The implication $4.\Rightarrow 1.$ is a direct consequence of Proposition 4.8. ∎
The next corollary shows that the conditions required by Proposition 4.9 can be relaxed if we do not require equality at every point, which is a condition that we do not need to prove our methods. Before we prove it, we need the following lemma.
In the next lemma, $\mathfrak{B}_{\mathbb{R}}$ denotes the Borel $\sigma$-algebra of $\mathbb{R}$.
Lemma 4.1.
Let $(\mathcal{X},\Sigma,\mathbb{P})$ be a probability space with $(\mathcal{X},\Sigma)$ being an analytic Borel space, and let $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ be a separable measurable space and $f:\mathcal{X}\to\mathcal{Y}$ a measurable function. If $\alpha:\mathcal{Y}\to\mathbb{R}$ is injective and $\Sigma_{\mathcal{Y}}=\alpha^{-1}(\mathfrak{B}_{\mathbb{R}})$, then there exists a measurable function $\beta:\mathbb{R}\to\mathcal{Y}$ such that $\beta\circ\alpha\circ f=f$ $\mathbb{P}$-a.e.
Proof.
Suppose that $\alpha:\mathcal{Y}\to\mathbb{R}$ is an injective function such that $\Sigma_{\mathcal{Y}}=\alpha^{-1}(\mathfrak{B}_{\mathbb{R}})$.
In the proof of Proposition 4.8, we demonstrated that our definition of analytic Borel spaces coincides with the definition of analytic spaces given by Cohn (2013). Therefore, using Corollary 8.4.3 and Lemma 8.6.1 of Cohn (2013), we can conclude that there exists $B\in\mathfrak{B}_{\mathbb{R}}$ such that $B\subseteq(\alpha\circ f)[\mathcal{X}]$ and $\mathbb{P}((\alpha\circ f)^{-1}[B])=1$.
Choose $y_{0}\in\mathcal{Y}$ and define $\beta:\mathbb{R}\to\mathcal{Y}$ by $\beta(t):=\begin{cases}\alpha^{-1}(t),&t\in B\\ y_{0},&t\notin B\end{cases}$.
Next, we will prove that $\beta$ is $(\mathfrak{B}_{\mathbb{R}},\Sigma_{\mathcal{Y}})$-measurable.
Let $E\in\Sigma_{\mathcal{Y}}$ be arbitrary. From the hypothesis, we know that there exists $A\in\mathfrak{B}_{\mathbb{R}}$ such that $E=\alpha^{-1}[A]$, allowing us to conclude that
$\beta^{-1}[E]=$ $(\beta^{-1}[E]\cap B)\cup(\beta^{-1}[E]\cap B^{c})=(\alpha[\alpha^{-1}[A]]\cap B)\cup(\beta^{-1}[E]\cap B^{c})$ (2)
$=$ $(A\cap\alpha[\mathcal{Y}]\cap B)\cup(\beta^{-1}[E]\cap B^{c})=(A\cap B)\cup(\beta^{-1}[E]\cap B^{c})$ (3)
Therefore, since $A,B\in\mathfrak{B}_{\mathbb{R}}$ and $\beta$ is constant on $B^{c}$, we infer that $\beta^{-1}[E]\in\mathfrak{B}_{\mathbb{R}}$, implying that $\beta$ is $(\mathfrak{B}_{\mathbb{R}},\Sigma_{\mathcal{Y}})$-measurable.
Since $(\alpha\circ f)(x)\in B$ for every $x\in(\alpha\circ f)^{-1}[B]$, we have $\beta(\alpha(f(x)))=f(x)$ for every $x\in(\alpha\circ f)^{-1}[B]$, implying that $\beta\circ\alpha\circ f=f$ $\mathbb{P}$-a.e., because $\mathbb{P}((\alpha\circ f)^{-1}[B])=1$.
Corollary 4.2.
Let $(\mathcal{X},\Sigma,\mathbb{P})$ be a probability space with $(\mathcal{X},\Sigma)$ being an analytic Borel space, and let $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ and $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ be separable measurable spaces. Suppose that $f:\mathcal{X}\to\mathcal{Y}$ and $g:\mathcal{X}\to\mathcal{Z}$ are measurable functions. If $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$, then there exists a measurable function $h:\mathcal{Y}\to\mathcal{Z}$ such that $g=h\circ f$ $\mathbb{P}$-a.e.
Proof.
Suppose that $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$.
By hypothesis, $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ is a separable measurable space; therefore, by Corollary D.7.3.III of Cavalcante (2026), there exists an injective measurable function $\alpha:\mathcal{Z}\to\mathbb{R}$ such that $\Sigma_{\mathcal{Z}}=\alpha^{-1}(\mathfrak{B}_{\mathbb{R}})$.
The hypothesis implies that $(\forall x,y\in\mathcal{X})(f(x)=f(y)\Rightarrow(\alpha\circ g)(x)=(\alpha\circ g)(y)\big)$, allowing us to conclude, using Proposition 4.9, that there exists a measurable function $h^{\prime}:\mathcal{Y}\to\mathbb{R}$ such that
$\alpha\circ g=h^{\prime}\circ f$ (4)
By Lemma 4.1, there exists a measurable function $\beta:\mathbb{R}\to\mathcal{Z}$ such that $\beta\circ\alpha\circ g=g$ $\mathbb{P}$-a.e. Using (4), this implies that $g=h\circ f$ $\mathbb{P}$-a.e., where $h:\mathcal{Y}\to\mathcal{Z}$ given by $h:=\beta\circ h^{\prime}$ is measurable. ∎
We now proceed to the proofs of the methods presented in Section 3.
Lemma 4.2.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model, $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $\mathcal{E}\ll\mu$, and $T:\mathcal{X}\to(\mathcal{T},\Sigma_{\mathcal{T}})$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space and $(\mathcal{T},\Sigma_{\mathcal{T}})$ is separable. For each $\theta\in\Theta$, choose $f_{\theta}\in[dP_{\theta}/d\mu]$. For any $x\in\mathcal{X}$ and $\Theta^{\prime}\subseteq\Theta$, define
$D(x,\Theta^{\prime}):=\big\{y\in\mathcal{X}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta^{\prime})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}.$
If there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ that is dense in $(\Theta,\tau\langle\mathcal{E}\rangle)$ such that, for any $x,y\in\mathcal{X}$ satisfying $T(x)=T(y)$, we have $y\in D(x,\Theta_{0})$, then $T$ is $\mathcal{E}$-sufficient.
Proof.
Suppose that there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ that is dense in $(\Theta,\tau\langle\mathcal{E}\rangle)$ such that, for any $x,y\in\mathcal{X}$ satisfying $T(x)=T(y)$, we have $y\in D(x,\Theta_{0})$.
Suppose that $\Theta_{0}=\{\theta_{n}\}_{n\in\mathbb{N}}$. From Proposition 4.7.i, we know that $P:\Sigma\to\mathbb{R}$ given by $P:=\sum_{n\in\mathbb{N}}2^{-(n+1)}P_{\theta_{n}}$ is a probability measure such that $\mathcal{E}\equiv P$.
By hypothesis, $\mathcal{E}\ll\mu$; therefore, $P\ll\mu$. Moreover, due to the almost everywhere uniqueness of a Radon-Nikodym derivative, we can conclude that $dP/d\mu=\sum_{n\in\mathbb{N}}2^{-(n+1)}f_{\theta_{n}}$ $\mu$-a.e.
Define $f:\mathcal{X}\to\overline{\mathbb{R}}$ by $f:=\sum_{n\in\mathbb{N}}2^{-(n+1)}f_{\theta_{n}}$. Then $f$ is a non-negative measurable function. Let $n\in\mathbb{N}$ be an arbitrary element. We have $\frac{dP_{\theta_{n}}}{d\mu}=\frac{dP_{\theta_{n}}}{dP}\frac{dP}{d\mu}$, which implies that $f_{\theta_{n}}=\frac{dP_{\theta_{n}}}{dP}f$ $\mu$-a.e.
Observe that $P(f\in\{0,\,\infty\})=0$, which implies that $P(f\in(0,\infty))=1$.
Define $g_{\theta_{n}}:\mathcal{X}\to\mathbb{R}$ by $g_{\theta_{n}}(x):=\begin{cases}\frac{f_{\theta_{n}}(x)}{f(x)},&f(x)\in(0,\infty)\\ 0,&f(x)\in\{0,\infty\}\end{cases}$.
Then $g_{\theta_{n}}$ is a non-negative measurable function. Moreover, we have $g_{\theta_{n}}=\frac{dP_{\theta_{n}}}{dP}$ $P$-a.e., which implies that $g_{\theta_{n}}\in[dP_{\theta_{n}}/dP]$.
The previous results hold for every $n\in\mathbb{N}$ since, by hypothesis, $n\in\mathbb{N}$ is arbitrary.
Define $S:\mathcal{X}\to\mathbb{R}^{\mathbb{N}}$ by $S(x)(n):=g_{\theta_{n}}(x)$. Then $S$ is measurable. Furthermore, it is straightforward to verify that $\sigma(S)=\vee_{n\in\mathbb{N}}\sigma(g_{\theta_{n}})$, allowing us to conclude, using Proposition 4.7.ii, that $\sigma(S)$ is $\mathcal{E}$-sufficient. Next, we will prove that, for any $x,y\in\mathcal{X}$ satisfying $T(x)=T(y)$, we have $S(x)=S(y)$.
Let $x,y\in\mathcal{X}$ be arbitrary elements such that $T(x)=T(y)$. Then, using the hypothesis, we obtain $y\in D(x,\Theta_{0})$, implying that there exists $h_{xy}\in(0,\infty)$ such that, for each $n\in\mathbb{N}$, we have $f_{\theta_{n}}(y)=f_{\theta_{n}}(x)h_{xy}$. Hence, using the definition of $f$, we conclude that $f(y)=f(x)h_{xy}$.
If $f(x)\in\{0,\infty\}$, then $f(y)\in\{0,\infty\}$, allowing us to conclude that $S(x)(n)=0=S(y)(n)$ for every $n\in\mathbb{N}$ and, consequently, that $S(x)=S(y)$. Suppose that $f(x)\notin\{0,\infty\}$. Then $f(y)\notin\{0,\infty\}$, which implies that for each $n\in\mathbb{N}$ we have
$S(y)(n)=\frac{f_{\theta_{n}}(y)}{f(y)}=\frac{f_{\theta_{n}}(x)h_{xy}}{f(x)h_{xy}}=\frac{f_{\theta_{n}}(x)}{f(x)}=S(x)(n)$
and, consequently, $S(x)=S(y)$.
Therefore, for any $x,y\in\mathcal{X}$ satisfying $T(x)=T(y)$, we have $S(x)=S(y)$, which implies, using Proposition 4.8 and the hypotheses that $(\mathcal{X},\Sigma)$ is analytic Borel and $(\mathcal{T},\Sigma_{\mathcal{T}})$ is separable, that $\sigma(S)\subseteq\sigma(T)$. Hence, recalling that we proved that $S$ is $\mathcal{E}$-sufficient, we conclude, using Corollary 8.5 in Heyer (1982), that $\sigma(T)$ is $\mathcal{E}$-sufficient and, consequently, that $T$ is $\mathcal{E}$-sufficient. ∎
The next corollary is an immediate consequence of Corollary 4.1 and Lemma 4.2, with $(\Theta,\tau\langle\mathcal{E}\rangle)$ replaced by $(\Theta,d_{\mathcal{E}})$.
Corollary 4.3.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model, $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $\mathcal{E}\ll\mu$, and $T:\mathcal{X}\to(\mathcal{T},\Sigma_{\mathcal{T}})$. Suppose that $(\mathcal{X},\Sigma)$ is an analytic Borel space and $(\mathcal{T},\Sigma_{\mathcal{T}})$ is separable. For each $\theta\in\Theta$, choose $f_{\theta}\in[dP_{\theta}/d\mu]$. For any $x\in\mathcal{X}$ and $\Theta^{\prime}\subseteq\Theta$, define $D(x,\Theta^{\prime}):=\big\{y\in\mathcal{X}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta^{\prime})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}$. If there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ that is dense in $(\Theta,d_{\mathcal{E}})$ such that, for any $x,y\in\mathcal{X}$ satisfying $T(x)=T(y)$, we have $y\in D(x,\Theta_{0})$, then $T$ is $\mathcal{E}$-sufficient.
Proof.
By Corollary 4.1, $\Theta_{0}$ dense in $(\Theta,d_{\mathcal{E}})$ implies $\Theta_{0}$ dense in $(\Theta,\tau\langle\mathcal{E}\rangle)$, so the conclusion follows directly from Lemma 4.2. ∎
Finally, we provide the proof of Method 3.1.
Proof of Method 3.1.
Suppose that $T$ is $\mathcal{E}$-sufficient and there exists a non-empty countable set $\Theta_{0}\subseteq\Theta$ such that, for any $x,y\in\mathcal{X}$ satisfying $y\in D(x,\Theta_{0})$, we have $T(x)=T(y)$.
From Proposition 4.2, we know that there exists a non-empty countable subset $\Theta_{1}\subseteq\Theta$ such that $\mathcal{E}\equiv\{P_{\theta}\}_{\theta\in\Theta_{1}}$. Define $\Theta_{2}:=\Theta_{0}\cup\Theta_{1}$. Then $\Theta_{2}$ is a non-empty countable subset of $\Theta$. Suppose $\Theta_{2}=\{\theta_{n}\}_{n\in\mathbb{N}}$. Denote by $\overline{\Theta_{2}}$ the closure of $\Theta_{2}$ in $\Theta$ w.r.t. the pseudo-metric $d_{\mathcal{E}}$. Then $\mathcal{E}\equiv\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$, since $\Theta_{1}\subseteq\overline{\Theta_{2}}$.
Consider the statistical model $\mathcal{F}:=((\mathcal{X},\Sigma),\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}})$. It is straightforward to verify that $\Theta_{2}$ is dense in $(\overline{\Theta_{2}},d_{\mathcal{F}})$ (because $d_{\mathcal{F}}$ is a restriction of $d_{\mathcal{E}}$), which implies, using Corollary 4.1, that $\Theta_{2}$ is dense in $(\overline{\Theta_{2}},\tau\langle\mathcal{F}\rangle)$. Therefore, using Proposition 4.7.i, we conclude that $P:\Sigma\to\mathbb{R}$ given by $P:=\sum_{n\in\mathbb{N}}2^{-(n+1)}P_{\theta_{n}}$ is a probability measure satisfying $\mathcal{F}\equiv P$.
By hypothesis, $\mathcal{E}\ll\mu$; therefore, $\mathcal{F}\ll\mu$, which implies $P\ll\mu$. Moreover, due to the almost everywhere uniqueness of a Radon-Nikodym derivative, we can conclude that $dP/d\mu=\sum_{n\in\mathbb{N}}2^{-(n+1)}f_{\theta_{n}}$ $\mu$-a.e.
Define $f:\mathcal{X}\to\overline{\mathbb{R}}$ by $f:=\sum_{n\in\mathbb{N}}2^{-(n+1)}f_{\theta_{n}}$. Then $f$ is a non-negative measurable function. Let $n\in\mathbb{N}$ be arbitrary. We have $\frac{dP_{\theta_{n}}}{d\mu}=\frac{dP_{\theta_{n}}}{dP}\frac{dP}{d\mu}$ $\mu$-a.e., which implies that $f_{\theta_{n}}=\frac{dP_{\theta_{n}}}{dP}f$ $\mu$-a.e.
Observe that $P(f\in\{0,\,\infty\})=0$, which implies $P(f\in(0,\infty))=1$.
Define $g_{\theta_{n}}:\mathcal{X}\to\mathbb{R}$ by $g_{\theta_{n}}(x):=\begin{cases}\frac{f_{\theta_{n}}(x)}{f(x)},&f(x)\in(0,\infty)\\ f_{\theta_{n}}(x),&f(x)\in\{0,\infty\}\end{cases}$. Then $g_{\theta_{n}}$ is a non-negative measurable function. Moreover, we have $g_{\theta_{n}}=\frac{dP_{\theta_{n}}}{dP}$ $P$-a.e., which implies that $g_{\theta_{n}}\in[dP_{\theta_{n}}/dP]$.
The previous results hold for every $n\in\mathbb{N}$ since, by hypothesis, $n\in\mathbb{N}$ is arbitrary.
Define $S:\mathcal{X}\to\mathbb{R}^{\mathbb{N}}$ by $S(x)(n):=g_{\theta_{n}}(x)$. Then $S$ is measurable. Furthermore, it is straightforward to prove that $\sigma(S)=\vee_{n\in\mathbb{N}}\sigma(g_{\theta_{n}})$, allowing us to conclude, using Propositions 4.6 and 4.7.ii, that $S$ is minimal $\mathcal{F}$-sufficient since $\{\theta_{n}\}_{n\in\mathbb{N}}$ is dense in $(\overline{\Theta_{2}},\tau\langle\mathcal{F}\rangle)$. Next, we will prove that, for any $x,y\in\mathcal{X}$ satisfying $S(x)=S(y)$, we have $T(x)=T(y)$.
Let $x,y\in\mathcal{X}$ be arbitrary elements satisfying $S(x)=S(y)$.
First, suppose that $f(x)\in\{0,\infty\}$. If $f(y)\notin\{0,\infty\}$, then
$f_{\theta_{n}}(x)=S(x)(n)=S(y)(n)=\frac{f_{\theta_{n}}(y)}{f(y)}$
for every $n\in\mathbb{N}$, which implies, recalling the definition of $f$, that
$f(x)=\sum_{n\in\mathbb{N}}2^{-(n+1)}f_{\theta_{n}}(x)=\frac{\sum_{n\in\mathbb{N}}2^{-(n+1)}f_{\theta_{n}}(y)}{f(y)}=1,$
contradicting the assumption $f(x)\in\{0,\infty\}$. Hence, $f(y)\in\{0,\infty\}$, which implies that, for each $n\in\mathbb{N}$, we have $f_{\theta_{n}}(x)=S(x)(n)=S(y)(n)=f_{\theta_{n}}(y)$ and, consequently, $y\in D(x,\Theta_{0})$ (because $\Theta_{0}\subseteq\Theta_{2}$). Therefore, using the hypothesis, we infer that $T(x)=T(y)$.
Now suppose that $f(x)\notin\{0,\infty\}$. Repeating the previous argument, we can conclude that $f(y)\notin\{0,\infty\}$, which implies that $\frac{f_{\theta_{n}}(x)}{f(x)}=S(x)(n)=S(y)(n)=\frac{f_{\theta_{n}}(y)}{f(y)}$ for each $n\in\mathbb{N}$, implying that $f_{\theta_{n}}(y)=f_{\theta_{n}}(x)h_{xy}$ for every $n\in\mathbb{N}$, where $h_{xy}:=\frac{f(y)}{f(x)}>0$. Therefore, we obtain $y\in D(x,\Theta_{0})$ (because $\Theta_{0}\subseteq\Theta_{2}$), which implies, using the hypothesis, that $T(x)=T(y)$.
From the previous cases, we conclude that $T(x)=T(y)$.
Therefore, for any $x,y\in\mathcal{X}$ satisfying $S(x)=S(y)$, we have $T(x)=T(y)$, which implies, using Corollary 4.2, that there exists a measurable function $h:\mathbb{R}^{\mathbb{N}}\to\mathcal{T}$ such that $T=h(S)$ $P$-a.e. Hence, $T=h(S)$ $P_{\theta}$-a.e. for every $\theta\in\overline{\Theta_{2}}$.
By hypothesis, $T$ is $\mathcal{E}$-sufficient; therefore, $T$ is $\mathcal{F}$-sufficient, allowing us to conclude, using the previous a.e. equality and the fact that $S$ is minimal $\mathcal{F}$-sufficient, that $T$ is minimal $\mathcal{F}$-sufficient. Hence, $T$ is $\mathcal{E}$-sufficient and minimal $\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$-sufficient, allowing us to conclude, using the definition of minimal sufficiency of statistics and recalling that $\mathcal{E}\equiv\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$ , that $T$ is minimal $\mathcal{E}$-sufficient. ∎
Before proving Method 3.2, we establish the following lemma.
Lemma 4.3.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $\mu:\Sigma\to\overline{\mathbb{R}}$ a $\sigma$-finite measure such that $\mathcal{E}\ll\mu$. For each $\theta\in\Theta$, choose $f_{\theta}\in[dP_{\theta}/d\mu]$. If $\Theta_{0}\subseteq\Theta$ is a subset such that, for each $\theta\in\Theta$, there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ satisfying $f_{\theta}=\lim_{n\to\infty}f_{\theta_{n}}$ $\mu$-a.e., then $\Theta_{0}$ is dense in $(\Theta,d_{\mathcal{E}})$.
Proof.
Suppose that $\Theta_{0}\subseteq\Theta$ is a subset such that, for each $\theta\in\Theta$, there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ satisfying $f_{\theta}=\lim_{n\to\infty}f_{\theta_{n}}$ $\mu$-a.e. Let $\theta\in\Theta$ and $\varepsilon>0$ be arbitrary elements. To conclude the proof, it suffices to prove that there exists $\theta^{\prime}\in\Theta_{0}$ such that $d_{\mathcal{E}}(\theta,\theta^{\prime})<\varepsilon$.
From the hypothesis, we know that there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ such that $f_{\theta}=\lim_{n\to\infty}f_{\theta_{n}}$ $\mu$-a.e. Since $\int_{\mathcal{X}}f_{\theta_{n}}d\mu=\int_{\mathcal{X}}f_{\theta}d\mu=1<\infty$ for every $n\in\mathbb{N}$, we can conclude, using Proposition 1.7.8 of Lerner (2014), that $\lim_{n\to\infty}\int_{\mathcal{X}}|f_{\theta_{n}}-f_{\theta}|d\mu=0$, which implies, using Proposition 4.3, that $\lim_{n\to\infty}d_{\mathcal{E}}(\theta_{n},\theta)=0$ and, consequently, that there exists $N\in\mathbb{N}$ such that $d_{\mathcal{E}}(\theta_{N},\theta)<\varepsilon$. ∎
Proof of Method 3.2.
Suppose that, for any $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff $y\in D(x)$.
Define $D(x,\Theta_{0}):=\big\{y\in\mathcal{X}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta_{0})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}$. Next, we will prove that, for any $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff $y\in D(x,\Theta_{0})$.
$(\Rightarrow)$ Suppose that $T(x)=T(y)$. Then, using the hypothesis, we conclude that $y\in D(x)$, which implies that $y\in D(x,\Theta_{0})$, since $D(x)\subseteq D(x,\Theta_{0})$. $\blacksquare$
$(\Leftarrow)$ Suppose that $y\in D(x,\Theta_{0})$. Then there exists $h_{xy}\in(0,\infty)$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\Theta_{0}$. Let $\theta\in\Theta$ be an arbitrary element. From the hypotheses, we know that there exists a sequence $(\theta_{n})_{n\in\mathbb{N}}$ of $\Theta_{0}$ such that the limit $\lim_{n\to\infty}f_{\theta_{n}}$ exists $\mu$-a.e. and for every $z\in\mathcal{X}$ we have
$f_{\theta}(z)=\begin{cases}\lim_{n\to\infty}f_{\theta_{n}}(z),&\text{if the limit exists,}\\ 0,&\text{otherwise}.\end{cases}$ (5)
Since $f_{\theta_{n}}(y)=f_{\theta_{n}}(x)h_{xy}$ for every $n\in\mathbb{N}$ and $h_{xy}\in(0,\infty)$, we can infer that $\lim_{n\to\infty}f_{\theta_{n}}(x)$ exists if and only if $\lim_{n\to\infty}f_{\theta_{n}}(y)$ exists, allowing us to conclude, using (5), that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$. As $\theta\in\Theta$ was arbitrary, it follows that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\Theta$, hence $y\in D(x)$. Therefore, by the hypothesis $T(x)=T(y)$. $\blacksquare$
From the previous implications, we conclude that, for any $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff $y\in D(x,\Theta_{0})$, which allows us to conclude, using Lemmas 4.2 and 4.3 and Method 3.1, that $T$ is minimal $\mathcal{E}$-sufficient. ∎
We conclude with the proof of Method 3.3 taken from Proposition 1.6.9 in Pfanzagl (1994).
Proof of Method $3.$
Suppose that, for any $a_{0},\cdots,a_{k}\in\mathbb{R}$ satisfying $\sum_{i=1}^{k}a_{i}\eta_{i}(\theta)=a_{0}$ for every $\theta\in\Theta$, we have $a_{i}=0$ for every $i\in\{0,\cdots,k\}$. Next, we will use Method 3.1 to obtain the desired result.
Define $T:\mathcal{X}\to\mathbb{R}^{k}$ by $T(x):=(T_{1}(x),\cdots,T_{k}(x))$. Then $T$ is measurable. Define $\eta:\Theta\to\mathbb{R}^{k}$ by $\eta(\theta):=(\eta_{1}(\theta),\cdots,\eta_{k}(\theta))$.
Denote by $\langle\cdot,\cdot\rangle$ the canonical inner product of $\mathbb{R}^{k}$. That is, $\langle a,b\rangle:=\sum_{i=1}^{k}a_{i}b_{i}$ for any $a:=(a_{1},\cdots,a_{k}),b:=(b_{1},\cdots,b_{k})\in\mathbb{R}^{k}$. From the hypothesis, we know that, for each $\theta\in\Theta$, we have $\frac{dP_{\theta}}{d\mu}(x)=e^{\langle\eta(\theta),T(x)\rangle-B(\theta)}h(x)$ $\mu$-a.e. Therefore, using the Neyman-Fisher Factorization Theorem, we conclude that $T$ is sufficient.
Define $\nu:\Sigma\to\overline{\mathbb{R}}$ by $\nu(E):=\int_{E}h(x)d\mu(x)$. Then $\nu$ is a $\sigma$-finite measure. Moreover, for any $E\in\Sigma$ and $\theta\in\Theta$, we have
$P_{\theta}(E)=\int_{E}e^{\langle\eta(\theta),T(x)\rangle-B(\theta)}h(x)d\mu(x)=\int_{E}e^{\langle\eta(\theta),T(x)\rangle-B(\theta)}d\nu(x),$
which allows us to conclude that $\mathcal{E}\ll\nu$ and
$\frac{dP_{\theta}}{d\nu}(x)=e^{\langle\eta(\theta),T(x)\rangle-B(\theta)}\,\,\nu\text{-a.e.}$ (6)
For each $\theta\in\Theta$, define $f_{\theta}:\mathcal{X}\to\mathbb{R}$ by $f_{\theta}(x):=e^{\langle\eta(\theta),T(x)\rangle-B(\theta)}$. From (6), we know that $f_{\theta}\in[dP_{\theta}/d\nu]$ for every $\theta\in\Theta$.
Denote by $\tau_{\mathbb{R}^{k}}$ the standard topology on $\mathbb{R}^{k}$.
For each subset $\Theta^{\prime}\subseteq\Theta$, define $\Delta(\Theta^{\prime}):=\{\eta(\theta):\theta\in\Theta^{\prime}\}$. Since $\Delta(\Theta)\subseteq\mathbb{R}^{k}$, we can conclude that $(\Delta(\Theta),\tau_{\mathbb{R}^{k}}|_{\Delta(\Theta)})$ is a separable metrizable space, which implies that there exists a non-empty countable subset $\Delta_{0}\subseteq\Delta(\Theta)$ that is dense in $(\Delta(\Theta),\tau_{\mathbb{R}^{k}}|_{\Delta(\Theta)})$, where $\tau_{\mathbb{R}^{k}}|_{\Delta(\Theta)}:=\{B\cap\Delta(\Theta):B\in\tau_{\mathbb{R}^{k}}\}$ is the subspace topology.
Since $\Delta_{0}\subseteq\Delta(\Theta)$, we can conclude that there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that $\Delta_{0}=\Delta(\Theta_{0})$.
Define $D(x,\Theta_{0}):=\big\{y\in\mathcal{X}:(\exists h_{xy}\in(0,\infty))(\forall\theta\in\Theta_{0})\big(f_{\theta}(y)=f_{\theta}(x)h_{xy}\big)\big\}$.
Let $x,y\in\mathcal{X}$ be arbitrary elements such that $y\in D(x,\Theta_{0})$. Next, we will prove that $T(x)=T(y)$. Then, there exists $h_{xy}\in(0,\infty)$ such that $f_{\theta}(y)=f_{\theta}(x)h_{xy}$ for every $\theta\in\Theta_{0}$, allowing us to infer, using the definition of $f_{\theta}$, that, for each $\theta\in\Theta_{0}$, we have the constant $e^{\langle\eta(\theta),T(y)-T(x)\rangle}=h_{xy}$ and, consequently, that
$\langle\eta(\theta),T(y)-T(x)\rangle=c$ (7)
for every $\theta\in\Theta_{0}$, where $c:=\log(h_{xy})$.
Since $\Delta(\Theta_{0})$ is dense in $(\Delta(\Theta),\tau_{\mathbb{R}^{k}}|_{\Delta(\Theta)})$, we can conclude, using (7) and the continuity of the function $\mathbb{R}^{k}\to\mathbb{R},\,t\mapsto\langle t,T(y)-T(x)\rangle$, that $\langle\eta(\theta),T(y)-T(x)\rangle=c$ for every $\theta\in\Theta$.
Therefore, using the definition of the inner product $\langle\cdot,\cdot\rangle$, we conclude that, for each $\theta\in\Theta$, we have $\sum_{i=1}^{k}\eta_{i}(\theta)(T_{i}(y)-T_{i}(x))=c$, which implies, using the hypothesis, that $T_{i}(y)-T_{i}(x)=0$ for every $i\in\{1,\cdots,k\}$ and, consequently, that $T_{i}(x)=T_{i}(y)$ for every $i\in\{1,\cdots,k\}$. Hence, we obtain $T(x)=T(y)$.
From the previous results, we conclude that $T$ is sufficient and that there exists a non-empty countable subset $\Theta_{0}\subseteq\Theta$ such that, for any $x,y\in\mathcal{X}$ satisfying $y\in D(x,\Theta_{0})$, we have $T(x)=T(y)$, which implies, using Method 3.1, that $T$ is minimal sufficient.
5 Acknowledgments
The first author acknowledges the financial support provided by CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior), Brazil, and by CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico), Brazil, Process No. 130535/2024-9.
Appendix A Measurable-Space Background
For convenience, we collect in this appendix the measurable-space background used in Sections 3 and 4.
Definition A.1.
A statistical model is a triple $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$, where $(\mathcal{X},\Sigma)$ is a measurable space, $\Theta$ is a non-empty set, and $\{P_{\theta}\}_{\theta\in\Theta}$ is a collection of probability measures on $\mathcal{X}$.
In practice, this model is often specified implicitly through a random sample, which is a measurable function $X:\Omega\to\mathcal{X}$ defined on a probability space such that the distribution induced by $X$ is $P_{\theta}$ for some $\theta\in\Theta$. A canonical example arises from a sequence of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},\cdots,X_{n}$, where each variable shares a common distribution law $Q_{\theta}$ for some $\theta\in\Theta$ and takes values in a measurable space $(\mathcal{Y},\Sigma_{\mathcal{Y}})$. This sequence represents the random sample $X:=(X_{1},\cdots,X_{n})$, whose probability law is the product measure $\otimes_{i=1}^{n}Q_{\theta}$, since $X_{1},\cdots,X_{n}$ are independent. Consequently, the statistical model associated with this random sample is $(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta}):=(\mathcal{Y}^{n},\otimes_{i=1}^{n}\Sigma_{\mathcal{Y}},\{\otimes_{i=1}^{n}Q_{\theta}\}_{\theta\in\Theta})$, where $\otimes_{i=1}^{n}\Sigma_{\mathcal{Y}}$ is the product $\sigma$-algebra. For brevity, a sequence of i.i.d. random variables $X_{1},\cdots,X_{n}$ itself is also commonly referred to as a random sample.
Definition A.2.
Let $\mathcal{E}:=(\mathcal{X},\Sigma,\{P_{\theta}\}_{\theta\in\Theta})$ be a statistical model and $(\mathcal{T},\Sigma_{\mathcal{T}})$ a measurable space. A statistic, i.e., a measurable function $T:\mathcal{X}\to\mathcal{T}$ is said to be ($\mathcal{E}$-)sufficient (w.r.t. that measurable space) iff for each $E\in\Sigma$, there exists a measurable function $\kappa_{E}:\mathcal{T}\to\mathbb{R}$ such that, for each $\theta\in\Theta$, we have $P_{\theta}(E|T)=\kappa_{E}(T)$ $P_{\theta}$-a.e. Furthermore, $T$ is said to be minimal ($\mathcal{E}$-)sufficient (w.r.t. measurable space) iff it is sufficient and, given any measurable space $(\mathcal{S},\Sigma_{\mathcal{S}})$ and sufficient statistic $S:\mathcal{X}\to(\mathcal{S},\Sigma_{\mathcal{S}})$, there exists a measurable function $f:\mathcal{S}\to\mathcal{T}$ such that, for each $\theta\in\Theta$, we have $T=f(S)$ $P_{\theta}$-a.e.
Informally, $T$ is sufficient iff the conditional probability $P_{\theta}(E|T)$ does not depend on $\theta$ for every $E\in\Sigma$ and $T$ is minimal sufficient iff it is sufficient and a function of any other sufficient statistic.
Definition A.3.
A standard Borel space is a measurable space whose $\sigma$-algebra is the Borel $\sigma$-algebra of a complete separable metric space.
Examples of standard Borel spaces include the Euclidean space $\mathbb{R}^{n}$, the extended real line $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$, and the space of all real sequences $\mathbb{R}^{\mathbb{N}}$, which is often denoted by $\mathbb{R}^{\infty}$. Furthermore, if $(\mathcal{X},\Sigma)$ is a standard Borel space and $B\in\Sigma$, then $(B,\Sigma|_{B})$ is also a standard Borel space, where $\Sigma|_{B}:=\{E\cap B:E\in\Sigma\}$ is the trace $\sigma$-algebra. This is a consequence of Corollary 13.4 in Kechris (1995).
Definition A.4.
A measurable space $(\mathcal{X},\Sigma)$ is said to be countably generated iff $\Sigma$ is generated by a countable collection of measurable sets.
Examples of such spaces include the Borel $\sigma$-algebras on second-countable topological spaces such as $\mathbb{R}^{n}$, $\overline{\mathbb{R}}$, and $\mathbb{R}^{\mathbb{N}}$. Moreover, every standard Borel space is countably generated. This is a consequence of Proposition 2.1.9 in Srivastava (1998).
Finally, we introduce two more concepts.
Definition A.5.
A measurable space $(\mathcal{X},\Sigma)$ is said to be separable iff it is countably generated and $\{x\}\in\Sigma$ for every $x\in\mathcal{X}$.
Every standard Borel space is separable since in a metric space every singleton is a closed set and, hence, measurable. Moreover, it can be proved that a measurable space is separable iff it is the Borel $\sigma$-algebra of a separable metric space, as can be seen in Proposition 4.3.10 of Doberkat (2015).
Definition A.6.
A measurable space $(\mathcal{X},\Sigma)$ is said to be analytic Borel iff it is separable and there exist a standard Borel space $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ and a surjective measurable function $f:\mathcal{Y}\to\mathcal{X}$.
Hence, every standard Borel space is an analytic Borel space.